# Gradient flow in parameter space is equivalent to linear interpolation in output space

### v1

@@ Abstract.
We prove that the usual gradient flow in parameter space that underlies many training algorithms for neural networks in deep learning can be continuously deformed into an adapted gradient flow which yields (constrained) Euclidean gradient flow in output space. Moreover, if the Jacobian of the outputs with respect to the parameters is full rank (for fixed training data), then the time variable can be reparametrized so that the resulting flow is simply linear interpolation, and a global minimum can be achieved.
@@ 1. Introduction
In [Che23], it was discussed in detail how the choice of the Riemannian structure in $\mathbb{R}_{\underline{\theta}}^{K}$ influences the training dynamics. In this context, a modified gradient flow that induces the Euclidean gradient flow in output space $\mathbb{R}_{\underline{x}}^{QN}$ was introduced and contrasted with the standard Euclidean gradient flow in parameter space; it was shown that both flows exhibit the same critical sets. In the work at hand, we extend these results by proving that those two flows are in fact homotopy equivalent to one another (Theorem 2.3). Moreover, we prove that if the Jacobian matrix exhibits no rank loss, then the Euclidean flow in output space is reparametrization equivalent to linear interpolation with respect to a suitable time variable (Proposition 2.4). When there is rank loss, an expression for the deviation from linear interpolation is given (Proposition 2.5).
@@
We briefly comment on related work in section 4. A short appendix on generalized inverses for rank deficient matrices is provided as well.
@@ 2. Main results
A common choice for cost function is the squared loss,
$\mathcal{C}(\underline{\theta},X_{0},Y)=\frac{1}{N}\left\lVert{f(\underline{\theta},X_{0})-Y}\right\rVert_{\mathcal{L}^{2}}^{2},$ (2.7)
where $\left\lVert{\cdot}\right\rVert_{\mathcal{L}^{2}}$ is the Frobenius or Hilbert-Schmidt norm.
@@
$\underline{y}:=(y(x_{1})^{T},\cdots,y(x_{N})^{T})^{T},$ (2.10)
In this case, the cost (2.7) becomes
$\mathcal{C}(\underline{x}(\underline{\theta}))=\frac{1}{N}\left\lvert{\underline{x}(\underline{\theta})-\underline{y}}\right\rvert^{2}.$ (2.11)
The Jacobian $D[\underline{\theta}]\in\mathbb{R}^{QN\times K}$ is
@@
$\underline{x}(s)$ $:=\underline{x}(\theta(s))$ (2.14)
@@
however, the two sets are not necessarily identical. It is known that if $\operatorname{rank}D[\underline{\theta}]=QN=\min\{K,QN\}$, then $\nabla_{\underline{x}}\mathcal{C}=0$ if, and only if, $\nabla_{\underline{\theta}}\mathcal{C}=0$. Moreover, if $\mathcal{C}$ is the squared loss, then $\nabla_{\underline{x}}\mathcal{C}=0$ corresponds to a global minimum with zero loss. See e.g. [Che23, CE23, KMTM24] for more details.
@@ 2.1. Adapted gradient flow
Let $K\geq QN$, and let $\underline{x}(s)=\underline{x}(\underline{\theta}(s))$ be defined as in (2.9). When $D$ is full rank, setting
@@
$\partial_{s}\underline{x}=-\mathcal{P}_{\operatorname{ran}(DD^{T})}\nabla_{\underline{x}}\mathcal{C}.$ (2.28)
@@
Let $\mathcal{C}:\mathbb{R}^{K}\to\mathbb{R}$ depend on $\underline{\theta}\in\mathbb{R}^{K}$ only through $\underline{x}(\underline{\theta})$, for $\underline{x}:\mathbb{R}^{K}\to\mathbb{R}^{QN}$ defined as in (2.9), and let $D[\underline{\theta}]\in\mathbb{R}^{QN\times K}$ be the Jacobian defined in (2.12). There is a one-parameter family of vector fields in $\mathbb{R}^{K}$ interpolating between
@@ 2.2. Reparametrization
If $\underline{x}(s)$ is a solution to Euclidean gradient flow (2.47), then

### v2

@@ Abstract.
We prove that the standard gradient flow in parameter space that underlies many training algorithms in deep learning can be continuously deformed into an adapted gradient flow which yields (constrained) Euclidean gradient flow in output space. Moreover, for the $L^{2}$ loss, if the Jacobian of the outputs with respect to the parameters is full rank (for fixed training data), then the time variable can be reparametrized so that the resulting flow is simply linear interpolation, and a global minimum can be achieved. For the cross-entropy loss, under the same rank condition and assuming the labels have positive components, we derive an explicit formula for the unique global minimum.
@@ 1. Introduction
In [Che23], it was discussed in detail how the choice of the Riemannian structure in $\mathbb{R}_{\underline{\theta}}^{K}$ influences the training dynamics. In this context, a modified gradient flow that induces the Euclidean gradient flow in output space $\mathbb{R}_{\underline{x}}^{QN}$ was introduced and contrasted with the standard Euclidean gradient flow in parameter space; it was shown that both flows exhibit the same critical sets. In the work at hand, we extend these results by proving that those two flows are in fact homotopy equivalent to one another (Theorem 2.3). Moreover, for the $L^{2}$ cost, we prove that if the Jacobian matrix exhibits no rank loss, then the Euclidean flow in output space is reparametrization equivalent to linear interpolation with respect to a suitable time variable (Proposition 2.4). When there is rank loss, an expression for the deviation from linear interpolation is given (Proposition 2.5). For the cross-entropy loss, we prove that output space fibers into invariant manifolds over which Euclidean gradient flow converges to a unique global minimum (Proposition 2.7).
@@
We briefly comment on related work in section 4. A short appendix on pseudoinverses for rank deficient matrices is provided as well.
@@ 2. Main results
The total cost $\mathcal{C}$ used for training is usually a sum of errors for each data point,
$\mathcal{C}(\underline{\theta})=\sum_{x\in\mathcal{X}_{0}}\ell(f(\underline{\theta},x),y(x)),$ (2.7)
where the chosen loss function $\ell:\mathbb{R}^{Q}\times\mathbb{R}^{Q}\to\mathbb{R}$ depends on the parameters only through $f(\underline{\theta},x).$
@@
$\underline{y}:=(y(x_{1})^{T},\cdots,y(x_{N})^{T})^{T}.$ (2.10)
In this case the cost can be redefined as $\mathcal{C}(\underline{x}(\underline{\theta}))$, and only depends on the parameters through its dependence on $\underline{x}(\underline{\theta})$. The Jacobian $D[\underline{\theta}]\in\mathbb{R}^{QN\times K}$ is
@@
$\underline{x}(s)$ $:=\underline{x}(\underline{\theta}(s))$ (2.13)
@@
however, the two sets are not necessarily identical. It is known that if $\operatorname{rank}D[\underline{\theta}]=QN=\min\{K,QN\}$, then $\nabla_{\underline{x}}\mathcal{C}=0$ if, and only if, $\nabla_{\underline{\theta}}\mathcal{C}=0$. Moreover, if $\mathcal{C}$ is the squared loss, then $\nabla_{\underline{x}}\mathcal{C}=0$ corresponds to a global minimum with zero loss. See e.g. [Coo18, TKB19, Che23, KMTM24, CE25] for more details.
Notation. Given a matrix $A\in\mathbb{R}^{m\times n}$, we denote by $(A){}^{+}$ its pseudoinverse, or Moore–Penrose inverse; see appendix A for a definition and useful properties. Moreover, given a subspace $V\subset\mathbb{R}^{d}$, we denote by $\mathcal{P}_{V}$ the orthogonal projection onto $V$, such that $\mathcal{P}_{V}=\mathcal{P}_{V}^{2}=\mathcal{P}_{V}^{T}$.
@@ 2.1. Adapted gradient flow
Let $K\geq QN$, and let $\underline{x}(s)=\underline{x}(\underline{\theta}(s))$ be defined as in (2.9). Assume $\underline{x}(\underline{\theta})$ has Lipschitz continuous derivatives. When $D$ is full rank, setting
@@
$\partial_{s}\underline{x}=-\mathcal{P}_{\operatorname{ran}(DD^{T})}\nabla_{\underline{x}}\mathcal{C}=-\mathcal{P}_{\operatorname{ran}(D)}\nabla_{\underline{x}}\mathcal{C}.$ (2.27)
@@
Let $\mathcal{C}:\mathbb{R}^{K}\to\mathbb{R}$ depend on $\underline{\theta}\in\mathbb{R}^{K}$ only through $\underline{x}(\underline{\theta})$, for $\underline{x}:\mathbb{R}^{K}\to\mathbb{R}^{QN}$ a function with Lipschitz continuous derivatives defined as in (2.9), and let $D[\underline{\theta}]\in\mathbb{R}^{QN\times K}$ be the Jacobian defined in (2.11). There is a one-parameter family of vector fields in $\mathbb{R}^{K}$ interpolating between
@@
This holds for any loss function $\ell$ in (2.7), and to any neural network whose activation function has Lipschitz continuous derivatives. Non-smooth activation functions, such as ReLU, can be mollified to satisfy this condition.
@@ 2.2. Reparametrization for the $L^{2}$ loss
If $\underline{x}(s)$ is a solution to Euclidean gradient flow (2.46) with respect to the cost (2.47), then
@@
2.3. Cross-entropy loss
For classification of data into $Q$ classes, the output $f_{\underline{\theta}}(x)$ and label $y(x)$ associated to an input $x\in\mathcal{X}_{0}$ are vectors in $\mathbb{R}^{Q}$ representing the probability that $x$ belongs to each class, so that $\ f_{\underline{\theta}}(x)_{j},y_{j}\in[0,1]$, $j=1,\cdots,Q$, and $\sum_{j=1}^{Q}f_{\underline{\theta}}(x)_{j}=\sum_{j=1}^{Q}y_{j}=1.$ The cross-entropy loss is
$\ell(f(x),y)=-y\cdot\log(f(x))=-\sum_{j=1}^{Q}y_{j}\log(f(x)_{j}),$ (2.64)
where the $\log$ is taken component-wise on vectors. In this case, the task becomes one of comparing the probability distribution of the true labels with that of the predicted labels. Prior work has explored the geometric properties of gradient flows on such probability manifolds within the framework of information geometry, see e.g. [Ama16, AJVLS17, LM18]. Whether the ideas presented here can be meaningfully extended to this setting remains an open question, which we leave for future investigation.
Let us consider instead a model whose output can be any vector in $\mathbb{R}^{Q}$. Let $\sigma$ be the softmax function, so that if $z\in\mathbb{R}^{Q}$, $\sigma(z)\in\mathbb{R}^{Q}$ and
$\sigma(z)_{j}=\frac{e^{z_{j}}}{\sum_{i}e^{z_{i}}},\quad j=1,\cdots,Q,$ (2.65)
and define
$\ell(f(x),y)=-\sum_{j=1}^{Q}y_{j}\log\sigma(f(x))_{j}=-y\cdot\log\sigma(f(x))$ (2.66)
with $y_{j}\in[0,1]$ for $j=1,\cdots,Q$ and $\sum_{i=1}^{Q}y_{i}=1$. Then Theorem 2.3 applies to
$\mathcal{C}(\underline{\theta})=\sum_{x\in\mathcal{X}_{0}}\ell(f_{\underline{\theta}}(x),y(x)).$ (2.67)
$\partial_{t}\underline{x}(t)$ $=-\nabla_{\underline{x}}\mathcal{C}(\underline{x}(t)).$ (2.68)
From now until the end of this subsection, we hide the dependence on the parameters $\underline{\theta}\in\mathbb{R}^{K}$ and let $f_{t}(x):=f(\underline{\theta}(t),x)$. We can compute, for some input $x\in\mathcal{X}_{0}$ and associated label $y=y(x)\in\mathbb{R}^{Q}$,
$\partial_{f(x)_{k}}\mathcal{C}$ $=\sum_{i\neq k}\left(y_{i}\frac{\sum_{j}e^{f(x)_{j}}}{e^{f(x)_{i}}}\frac{e^{f(x)_{i}}e^{f(x)_{k}}}{\left(\sum_{j}e^{f(x)_{j}}\right)^{2}}\right)-y_{k}\frac{\sum_{j}e^{f(x)_{j}}}{e^{f(x)_{k}}}\frac{\left(\sum_{j}e^{f(x)_{j}}\right)e^{f(x)_{k}}-e^{2f(x)_{k}}}{\left(\sum_{j}e^{f(x)_{j}}\right)^{2}}$
$=\sum_{i\neq k}\left(y_{i}\frac{e^{f(x)_{k}}}{\sum_{j}e^{f(x)_{j}}}\right)-y_{k}\frac{\sum_{j}e^{f(x)_{j}}-e^{f(x)_{k}}}{\sum_{j}e^{f(x)_{j}}}$
$=-y_{k}+\frac{e^{f(x)_{k}}}{\sum_{j}e^{f(x)_{j}}}\sum_{i=1}^{Q}y_{i}$ (2.69)
$=-y_{k}+\frac{e^{f(x)_{k}}}{\sum_{j}e^{f(x)_{j}}},$
and end up with a system of ordinary differential equations for each input:
$(f_{t}(x)_{j})^{\prime}=-\partial_{f_{t}(x)_{j}}\mathcal{C}=y_{j}-\frac{e^{f_{t}(x)_{j}}}{\sum_{i}e^{f_{t}(x)_{i}}},\quad j=1,\cdots,Q,$ (2.70)
or
$f_{t}(x)^{\prime}=-\nabla_{f_{t}(x)}\mathcal{C}=y-\sigma(f_{t}(x)).$ (2.71)
We make two key observations: First, note that
$\sum_{j=1}^{Q}f_{t}(x)_{j}^{\prime}=\underbrace{\sum_{j=1}^{Q}y_{j}}_{=1}-\underbrace{\frac{\sum_{j=1}^{Q}e^{f_{t}(x)_{j}}}{\sum_{i=1}^{Q}e^{f_{t}(x)_{i}}}}_{=1}=0,$ (2.72)
so
$\sum_{j=1}^{Q}f_{t}(x)_{j}(s)=\sum_{j=1}^{Q}f_{0}(x)_{j}=:c.$ (2.73)
Thus, even though we assumed that $f_{t}(x)\in\mathbb{R}^{Q}$ is not necessarily on the probability simplex, so that $c\in\mathbb{R}$ can have any value, a solution to (2.71) is still constrained to an affine hyperplane
$H_{c}:=\left\{z\in\mathbb{R}^{Q}:z\cdot u_{Q}=c\right\}$ (2.74)
where $u_{Q}:=(1,\cdots,1)^{T}\in\mathbb{R}^{Q}$.
Second, we compute the Hessian of the loss (2.66) for a single data point,
$\left[D^{2}\ell\right]_{jk}$ $=\partial_{f_{t}(x)_{k}}\left(-y_{j}+\frac{e^{f_{t}(x)_{j}}}{\sum_{i}e^{f_{t}(x)_{i}}}\right)$
$=\delta_{jk}\frac{e^{f_{t}(x)_{j}}}{\sum_{i}e^{f_{t}(x)_{i}}}-\frac{e^{f_{t}(x)_{j}}}{\sum_{i}e^{f_{t}(x)_{i}}}\frac{e^{f_{t}(x)_{k}}}{\sum_{i}e^{f_{t}(x)_{i}}},$ (2.75)
so
$D^{2}\ell\,\rvert_{f_{t}(x)}=\operatorname{diag}(\sigma(f_{t}(x)))-\sigma(f_{t}(x))\sigma(f_{t}(x))^{T}.$ (2.76)
Observe that
[…]
Lemma 2.6.
Let $D^{2}\ell\,\rvert_{z}\in\mathbb{R}^{Q\times Q}$ be the Hessian matrix of the loss (2.66) evaluated at an arbitrary point $z\in\mathbb{R}^{Q}$,
$D^{2}\ell\,\rvert_{z}=\operatorname{diag}(\sigma(z))-\sigma(z)\sigma(z)^{T}.$ (2.78)
For all $z\in\mathbb{R}^{Q}$, $\operatorname{rank}D^{2}\ell\,\rvert_{z}=Q-1$ and $D^{2}\ell\,\rvert_{z}$ is positive-semidefinite.
Proof.
Fix $z\in\mathbb{R}^{Q}$. First, since $\sigma(z)_{j}>0$ for all $j=1,\cdots,Q$, we have $\operatorname{rank}\operatorname{diag}(\sigma(z))=Q$. Moreover, $\operatorname{rank}(\sigma(z)\sigma(z)^{T})=1$, so
$\operatorname{rank}D^{2}\ell\,\rvert_{z}\geq\operatorname{rank}\operatorname{diag}(\sigma(z))-\operatorname{rank}(\sigma(z)\sigma(z)^{T})=Q-1.$ (2.79)
On the other hand,
$D^{2}\ell\,\rvert_{z}u_{Q}=0,$ (2.80)
so $\dim(\ker D^{2}\ell\,\rvert_{z})\geq 1$. Therefore,
$Q-1\leq\operatorname{rank}D^{2}\ell\,\rvert_{z}=Q-\dim(\ker D^{2}\ell\,\rvert_{z})\leq Q-1.$ (2.81)
To show $D^{2}\ell\,\rvert_{z}\geq 0$, let $S:=\operatorname{diag}(\sigma(z))$. Then for any $v\in\mathbb{R}^{Q}$,
$\langle v,D^{2}\ell\,\rvert_{z}v\rangle$ $=\langle v,Sv\rangle-\left\lvert{\langle v,\sigma(z)\rangle}\right\rvert^{2}.$ (2.82)
Using Cauchy–Schwarz,
$\left\lvert{\langle v,\sigma(z)\rangle}\right\rvert^{2}$ $=\left\lvert{\langle S^{\frac{1}{2}}v,S^{-\frac{1}{2}}\sigma(z)\rangle}\right\rvert^{2}$ (2.83)
$\leq\langle v,Sv\rangle\langle\sigma(z),S^{-1}\sigma(z)\rangle.$ (2.84)
Therefore
$\langle v,D^{2}\ell\,\rvert_{z}v\rangle\geq\langle v,Sv\rangle\left(1-\langle\sigma(z),S^{-1}\sigma(z)\rangle\right),$ (2.85)
but
$S^{-1}\sigma(z)=u_{Q}$ (2.86)
and
$\langle\sigma(z),S^{-1}\sigma(z)\rangle=\sigma(z)\cdot u_{Q}=\sum_{j=1}^{Q}\sigma(z)_{j}=1,$ (2.87)
thus
$\langle v,D^{2}\ell\,\rvert_{z}v\rangle\geq 0.$ (2.88)
From this we conclude that $D^{2}\ell\,\rvert_{z}$ is positive-definite when restricted to the submanifolds $H_{c}$, so that for each training data point the cross-entropy loss (2.66) appears convex to a solution of the Euclidean gradient flow (2.71). If there exists a critical point, it should correspond to a global minimum. Such an equilibrium point must satisfy
$\sigma(f_{*}(x))=y,$ (2.89)
and this can only be attained if $y_{j}>0$, for all $j=1,\cdots,Q$111 While it is common to train classifiers using one-hot encoded labels, where $y\in\mathbb{R}^{Q}$ are canonical basis vectors, training with labels that assign positive probabilities to all classes leads to better convergence properties, as we show in the present work. This scenario can occur when certain training inputs do not clearly belong to a single class, or can be forced by subjecting the labels to small perturbations. We have thus proved most of the following:
Proposition 2.7.
If $f_{t}(x)$ is a solution to the gradient flow (2.71) with respect to the cost (2.67), then its image is restricted to an affine hyperplane,
$\sum_{j=1}^{Q}f_{t}(x)_{j}=c,$ (2.90)
where $c=\sum_{j=1}^{Q}f_{0}(x)_{j}$.
Moreover, if the associated label $y\in\mathbb{R}^{Q}$ has positive components, then $f_{t}(x)\to f_{*,c}(x)$ as $t\to\infty$, with
$f_{*,c}(x)=\log y+\frac{1}{Q}\left(\sum_{j=1}^{Q}(f_{0}(x)_{j}-\log y_{j})\right)u_{Q},$ (2.91)
where $\log y=(\log y_{1},\cdots,\log y_{Q})^{T}$.
Note that
$\nabla_{\underline{x}(t)}\mathcal{C}=(\nabla_{f_{t}(x_{1})}\ell(f_{t}(x_{1}),y(x_{1}))^{T},\cdots,\nabla_{f_{t}(x_{N})}\ell(f_{t}(x_{N}),y(x_{N}))^{T})^{T},$ (2.92)
and as with the $L^{2}$ loss, the dynamics of training for each data point in the set of training data $\mathcal{X}_{0}$ is independent. Further, the Hessian of the total cost (2.67) is block-diagonal,
$D^{2}\mathcal{C}\,\rvert_{\underline{x}}=\operatorname{diag}(D^{2}\ell\rvert_{f_{t}(x_{1})},\cdots,D^{2}\ell\rvert_{f_{t}(x_{N})})\in\mathbb{R}^{QN\times QN}.$ (2.93)
It is straightforward to check that $D^{2}\mathcal{C}\,\rvert_{\underline{x}}$ is positive-definite when restricted to a product of hyperplanes $H_{c_{1}}\times\cdots\times H_{c_{N}}\subset\mathbb{R}^{QN}$. Therefore, we may extend Proposition 2.7 accordingly.
Corollary 2.8.
If $\underline{x}(t)=(f_{t}(x_{1})^{T},\cdots,f_{t}(x_{N})^{T})^{T}\in\mathbb{R}^{QN}$ is a solution to Euclidean gradient flow (2.68) with respect to the cost (2.67), then $\underline{x}(t)$ is constrained to a product of $N$ affine hyperplanes in $\mathbb{R}^{Q}$, defined by
$\sum_{j=1}^{Q}f_{t}(x_{n})_{j}=c_{n},$ (2.94)
where $c_{n}=\sum_{j=1}^{Q}f_{0}(x_{n})_{j}$, for $n=1,\cdots,N$.
Moreover, if the associated label vector $\underline{y}\in\mathbb{R}^{QN}$ has positive components, then there exists a family of global minima $\underline{x}^{*}(C)$ parametrized by $C=(c_{1},\cdots,c_{N})\in\mathbb{R}^{N}$, and $\underline{x}(t)$ converges to a unique equilibrium
$\underline{x}^{*}(C)$ $=(f_{*,c_{1}}(x_{1})^{T},\cdots,f_{*,c_{N}}(x_{N})^{T})^{T}$ (2.95)
as $t\to\infty$, for $f_{*,c}$ defined in (2.91).
Proof.
All that is left to prove Proposition 2.7 and Corollary 2.8 is to find a formula for $f_{*,c}(x)$. Suppose that $f_{t}(x)$ is such that
$e^{f_{t}(x)_{j}}=\eta y_{j},\quad j=1,\cdots,Q,$ (2.96)
so that
$\sigma(f_{t}(x))_{j}=\frac{e^{f_{t}(x)_{j}}}{\sum_{i}e^{f_{t}(x)_{i}}}=\frac{\eta y_{j}}{\sum_{i}\eta y_{i}}=y_{j}.$ (2.97)
This satisfies $f_{t}(x)^{\prime}=0$ for (2.71), and is thus an equilibrium point for the gradient flow. Then
$f_{*,c}(x)_{j}=\log(\eta y_{j})=\log(y_{j})+\log(\eta),\quad j=1,\cdots,Q.$ (2.98)
If
$\sum_{j=1}^{Q}f_{0}(x)_{j}=c,$ (2.99)
then
$c=\sum_{j=1}^{Q}f_{*,c}(x)_{j}=\sum_{j=1}^{Q}\log y_{j}+Q\log(\eta)$ (2.100)
and
$\log(\eta)=\frac{1}{Q}\left(c-\sum_{j}\log y_{j}\right),$ (2.101)
from which we conclude
$f_{*,c}(x)_{j}$ $=\log y_{j}+\frac{1}{Q}\left(c-\sum_{i=1}^{Q}\log y_{i}\right)$
$=\log y_{j}+\frac{1}{Q}\sum_{i=1}^{Q}(f_{0}(x)_{i}-\log y_{i})$ (2.102)
for each $j=1,\cdots,Q.$ Since $D^{2}\ell\,\rvert_{f_{t}(x)}$ is positive-definite on the hyperplane $H_{c}$, this minimum is unique on $H_{c}$.
Finally, as
$\ell(f_{t}(x),y)=-y\cdot\log\sigma(f_{t}(x)),$ (2.103)
we have
$\ell(f_{*,c}(x),y)=-y\cdot\log y>0$ (2.104)
which does not depend on the initial value $f_{0}(x)$. Therefore, $f_{*,c}(x)$ is a degenerate global minimum. ∎
Remark 2.9.
For a fixed $x\in\mathcal{X}_{0}$ and two different sets of initial conditions,
$f_{*,c_{1}}(x)-f_{*,c_{2}}(x)=\frac{1}{Q}(c_{2}-c_{1})u_{Q}$ (2.105)
or
$f_{*,c_{1}}(x)=f_{*,c_{2}}(x)+\frac{c_{2}-c_{1}}{Q}u_{Q}.$ (2.106)
This implies that the equilibrium manifold is a straight line passing through $f_{*,c}(x)$, for any fixed $c\in\mathbb{R}$. We conclude that $\mathbb{R}^{Q}$ fibers into invariant manifolds $H_{c}$ for the adapted Euclidean gradient flow, and that the critical manifold $Crit(x)=\left\{z:\nabla_{z}\ell(z,y(x))=0\right\}$ is given by $\left\{f_{*,c}(x)+\lambda u_{Q}:c\in\mathbb{R}\text{ fixed},\lambda\in\mathbb{R}\right\}$. Moreover, $Crit(x)$ and $H_{c}$ intersect transversally, for all $c\in\mathbb{R}.$